4.) True or False : 3Z = Z + Z + Z when Z is a matrix.

􏰀1 2􏰁 􏰀a􏰁

5.) Let X = 3 4 ; E = b

Find each of the following. If the operation cannot be done : state undefined operation.

a) XE

b) EX

c) XT X where XT stands for the transpose of X


10.) Consider the linear equation 2a + 3b = 4

Is (a; b) = ( 1/2 ; 1) a solution to the equation? Motivate your answer.


11.) Look up what is meant by a system of linear equations.

A known fact of solutions of systems of linear equations is that only one the following options can hold :

(a) No solution possible

(b) A unique solution can be found

(c) The system has infinite solutions.

Consider that two straight lines form a linear system.

Interpret what happens geometrically to the straight lines to get each case of the solution types given above.


12.) Look up the concept of a homogeneous linear system.

Only two solution types of the three mentioned solution types above are possible. Which one can never happen and why.


1
Expert's answer
2021-10-07T15:32:59-0400

4)

Let


"Z=\begin{pmatrix}
 a_{11} & a_{12} & ... & a_{1m} \\
 a_{21} & a_{22} & ... & a_{2m} \\
 \vdots & \vdots & \vdots & \vdots \\
 a_{n1} & a_{n2} & ... & a_{nm} \\
\end{pmatrix}"

Then


"3Z=\begin{pmatrix}
 3a_{11} & 3a_{12} & ... & 3a_{1m} \\
 3a_{21} & 3a_{22} & ... & 3a_{2m} \\
 \vdots & \vdots & \vdots & \vdots \\
 3a_{n1} & 3a_{n2} & ... & 3a_{nm} \\
\end{pmatrix}"

"Z+Z+Z=\begin{pmatrix}
 a_{11} & a_{12} & ... & a_{1m} \\
 a_{21} & a_{22} & ... & a_{2m} \\
 \vdots & \vdots & \vdots & \vdots \\
 a_{n1} & a_{n2} & ... & a_{nm} \\
\end{pmatrix}"

"+\begin{pmatrix}
 a_{11} & a_{12} & ... & a_{1m} \\
 a_{21} & a_{22} & ... & a_{2m} \\
 \vdots & \vdots & \vdots & \vdots \\
 a_{n1} & a_{n2} & ... & a_{nm} \\
\end{pmatrix}+\begin{pmatrix}
 a_{11} & a_{12} & ... & a_{1m} \\
 a_{21} & a_{22} & ... & a_{2m} \\
 \vdots & \vdots & \vdots & \vdots \\
 a_{n1} & a_{n2} & ... & a_{nm} \\
\end{pmatrix}"

"=\begin{pmatrix}
 a_{11}+a_{11}+a_{11} & a_{12}+a_{12}+a_{12} & ... & a_{1m}+a_{1m}+a_{1m} \\
 a_{21}+a_{21}+a_{21} & a_{22}+a_{22}+a_{22} & ... & a_{2m} +a_{2m}+a_{2m}\\
 \vdots & \vdots & \vdots & \vdots \\
 a_{n1}+a_{n1}+a_{n1} & a_{n2}+a_{n2}+a_{n2} & ... & a_{nm}+a_{nm}+a_{nm} \\
\end{pmatrix}"


"=\begin{pmatrix}
 3a_{11} & 3a_{12} & ... & 3a_{1m} \\
 3a_{21} & 3a_{22} & ... & 3a_{2m} \\
 \vdots & \vdots & \vdots & \vdots \\
 3a_{n1} & 3a_{n2} & ... & 3a_{nm} \\
\end{pmatrix}=3Z"

"3Z = Z + Z + Z" when "Z" is a matrix is True.


5)


"X=\begin{pmatrix}
 1 & 2 \\
 3 & 4
\end{pmatrix}, E=\begin{pmatrix}
 a \\
 b
\end{pmatrix}"

a)

"XE=\begin{pmatrix}
 1 & 2 \\
 3 & 4
\end{pmatrix}\begin{pmatrix}
 a \\
 b
\end{pmatrix}=\begin{pmatrix}
 a +2b \\
 3a+4b
\end{pmatrix}"

b)

The matrix "E" is "2\times1" matrix, the matrix "X" is "2\times2" matrix.

Since "1\not=2," then


"EX=\begin{pmatrix}
 a \\
 b
\end{pmatrix}\begin{pmatrix}
 1 & 2 \\
 3 & 4
\end{pmatrix}=does\ not\ exist"

c)


"X^T=\begin{pmatrix}
 1 & 3 \\
 2 & 4
\end{pmatrix}"

"X^TX=\begin{pmatrix}
 1 & 3 \\
 2 & 4
\end{pmatrix}\begin{pmatrix}
 1 & 2 \\
 3 & 4
\end{pmatrix}=\begin{pmatrix}
 1+9 & 2+12 \\
 2+12 & 4+16
\end{pmatrix}"

"=\begin{pmatrix}
 10 & 14 \\
 14 & 20
\end{pmatrix}"

10.) Consider the linear equation

"2a + 3b = 4"

If "(a, b)=(\dfrac{1}{2}, 1)," then substitute


"2(\dfrac{1}{2}) + 3(1) = 4"

"4=4, True"

Therefore "(a, b)=(\dfrac{1}{2}, 1)" is a solution to the equation "2a+3b=1."


11)

a) Two lines are parallel lines or skew lines.


b) Two lines are intersecting lines.


c) Two lines are coincident lines.


12) Since a homogeneous system always has a solution (the trivial solution), it can never be inconsistent.

Therefore a homogeneous linear system can have:

(a) A unique solution.

Or

(c) Infinite solutions.